# Non-Prehensile Tool-Object Manipulation by Integrating LLM-Based Planning and Manoeuvrability-Driven Controls

### v5

@@ Abstract
[…]
t1t1footnotetext: This work is supported in part by the Research Grants Council of Hong Kong under grant 15212721 and in part by the Jiangsu Industrial Technology Research Institute Collaborative Research Program Scheme under grant ZG9V.
@@ 1.1 Related Works
Transitioning from the critical role of task planning, it is evident that effective tool use is inherently tied to understanding the relationship between tools and objects [31]. Indeed, the success of a given tool-object manipulation task largely depends on the appropriate selection of the tool, which necessitates a nuanced comprehension of how tools interact with various objects in their environment. For example, robots can identify the tool type, potential uses, and contact approaches based on the tool’s geometry, see e.g., [14, 2]. In [32], tool features are learned through observation of the task’s effects and experimental validation of feature hypotheses. Affordance models are a common technique used for tool feature selection [33, 34, 35] and tool classification [36, 35, 37]. The relation between tool actions and its effects on objects is explored in [38, 37], where robots acquire affordance knowledge through predefined actions (e.g., pull, push, rotate). Recently, researchers have also explored the use of LLM in accelerating affordance learning in tool manipulation [2]. Some works have studied tool-based manipulation under constraints and from demonstrations [39]. Non-prehensile object manipulation strategies have been used in [40, 41].
@@ 1.2 Contributions
To address this research gap, in this work, we propose a novel LLM-based manoeuvrability-driven method with the following original contributions: (1) We develop an effective model to represent the geometric-mechanical relations and manoeuvrability of tools and objects; (2) We propose a non-prehensile strategy to manoeuvre objects under different constraints with tools; (3) We evaluate the performance of the proposed methodology with real-world experiments on a dual-arm robotic system. Our work uniquely integrates LLMs to enhance tool-object interactions, enabling robots to interpret and execute complex non-prehensile manipulation tasks through natural language instructions. This integration allows for dynamic adaptation to various situations and fosters intuitive human-robot collaboration, significantly improving the effectiveness of dual-arm tool-object manipulation.
The rest of the manuscript is organised as follows: Sec. 2 presents the methodology, Sec. 3 presents the results, Sec. 4 gives final conclusions.
@@ 2.1 Problem Formulation
We define a tool as a manipulable object that is graspable by a robot, a manipulandum [14] as an object (e.g. a block) that is manipulated via a tool, and a wall as a static non-manipulable object. Tool use by robots is challenging as the tools can have various shapes, the environment can be dynamic, and the contact between the tool and the manipulandum may be hard to maintain in a long-horizon task. In this study, we focus on using the side part of a tool to interact with the manipulandum. Depending on the geometric features of a tool and a wall, the available affordance for manoeuvring a manipulandum may be different. Affordance here refers to the available action-effects offered by the tool or the environment. In this work, we classify affordance into two types: active and passive. Active affordance is given from a manipulable object, i.e. a tool, and it is directly related to the manoeuvrability when driving a manipulandum. Passive affordance is given from a static non-manipulable object.
@@ 2.2 LLM-Based High-Level Symbolic Task Planner
To obtain a valid task decomposition for a long-horizon task, the system needs to understand the requirements and generate an executable subtask list. We develop a symbolic task planner that takes natural language instructions with scene descriptions as input, and outputs a list of high-level subtasks. The list involves the tool selection/sharing between two arms, the sequence to manipulate the tool with the manipulandum, and the interaction between the two arms. The model is fine-tuned using approximately 20,000 example data lists, specifically tailored for our non-prehensile tool object manipulation scenario. During the fine-tuning stage, we utilized a program to create 20,000 distinct environmental setups by randomly varying the poses of the robot, tool, block, and target within a finite combination space. This approach allows task decomposition to be framed as a classification problem, enabling the LLM to effectively correlate each setup with a specific list of motion functions, thereby enhancing its ability to predict expected outcomes based on prompt patterns rather than producing hallucinations.
@@
The LLM interprets the output of $f(\mathbf{L},\mathbf{o})$ to generate a subtask list $\{\mathbf{l}_{1},\mathbf{l}_{2},\dots\}\leftarrow f_{\text{llm}}(f(\mathbf{L},\mathbf{o}))$ where $\mathbf{l}_{i}$ is a subtask describing the manipulation phase of each robot and is corresponding to a high-level robot motion function. The motion functions are designed to be simple and specify a short-term goal of the concerned object (these functions omit low-level motion commands). For simplicity, here we use m to represent manipulandum in the following function definitions. We use grasp(arm, tool) for grasping a tool with the robot arm; approach(arm, tool, m) for approaching the location of m with tool using arm; interact(arm, tool, m, goal) for moving m to the goal location with the tool; stepping(arm, tool, m) for moving m out from the bounded area with the tool of the arm through contact pulsing motions; pass(arm1, tool, m, arm2) for passing m to another arm’s workspace; release(arm, tool) for releasing the tool back to its original place with the arm.
@@ 2.4 Manoeuvrability Analysis
A tool can push the manipulandum from the side, from the tip, or other areas. However, the relative location of the manipulandum respective to the tool affects its manoeuvrability. In other words, the affordance provided by the tool is proportional to manoeuvrability. Consider using a rotating stick to push an object with its end tip. In this situation, the tool may lose contact with the manipulandum as it rolls outwards, hence, the manoeuvrability of this point is low. On the other hand, the midpoint of the stick has a high manoeuvrability, which proportionally decreases as the contact point is further away from the midpoint. This behaviour can be modelled with a Gaussian function, where its centre is the segment’s centre and the peak height is half the segment’s length, see Fig. 3(b). We refer to this region as an affordance area.
@@
Tool regions with high values in the image $\mathbf{M}$ reflect a high manoeuvrability. These computed manoeuvrability values are useful to determine the location where the tool interacts with the manipulandum. To determine the centre of the object, we then expand the contour of the tool by the object’s radius $r^{\text{obj}}$. This contour is downsamppled with the Ramer-Douglas-Peucker algorithm, then, parameterised with the spline fitting technique. To extract key features of the tool geometry, we use a sliding window strategy to examine a small number of neighbouring points. Let $\mathcal{C}$ be the contour of the tool expanded by $r^{\text{obj}}$. The key features of the tool geometry are extracted using the following equation:
@@
To compute the minimal number of key points (denoted as $\mathbf{P}^{\text{key}}=\{\mathbf{p}^{\text{key}}_{1},\mathbf{p}^{\text{key}}_{2},\dots\}$) that capture the highest manoeuvrability among feature points, we use the density-based clustering algorithm. By integrating the affordance areas we obtained earlier, we can filter out some redundant key points. For example, if there exists a point $\mathbf{p}^{\text{key}}_{i}$ located outside the affordance area (visualised in Fig. 3(b)), we consider this point as redundant. All the non-redundant points are then grouped into $\mathbf{P}^{\diamond}=\{\mathbf{p}^{\diamond}_{1},\mathbf{p}^{\diamond}_{2},\dots\}$. To find the point in $\mathbf{P}^{\diamond}$ with the highest manoeuvrability (defined as $\mathbf{p}_{\ast}$), we use the manoeuvrability matrix $\mathbf{M}$ and distance between $\mathbf{p}^{\diamond}_{i}$ and $\mathbf{a}_{\ast}$ as described in the metric below:
@@ 2.5.2 Motion Strategy
where the coordinates of $\mathbf{p}^{\text{tool}}$ can be determined by finding a point $\mathbf{p}=(x,y)$ where it minimizes the distance between $(\mathbf{p}^{*},\mathbf{p}^{\text{obj}})$ with $f(\mathbf{p})$ and $(\mathbf{p}^{\text{tool}},\mathbf{p}^{\text{int}}_{\text{start}})$; then translates along the $x$ and $y$ axes until it reaches $\mathbf{p}^{\text{int,x,y}}_{\text{end}}$ with $k_{\text{int}}(\mathbf{p}^{\text{int,x,y}}_{\text{end}}-\mathbf{p}^{\text{tool}})$, where $k_{\text{int}}$ is determined empirically; lastly, the tool is rotated to align with the orientation of $\mathbf{p}^{\text{int}}_{\text{end}}$.
@@ 2.6 Application with Environmental Constraints
Consider the manipulandum is tightly confined within a concave-shaped wall, as shown in Fig. 5(a), with an unknown exit and assume that the tool can enter the constrained area. To move the manipulandum out from the bounded area with small movement space, we determine the direction from the manipulandum to the exit by considering the overall affordance of the wall boundary. We denote this direction vector as $\mathbf{v}^{\text{exit}}$, and its magnitude is defined as the minimum travel distance for the manipulandum. Consider the inner edge of the wall as a segment $\mathbf{s}^{\text{wall}}_{i}$ where $i=\{1,\dots,n^{\text{wall}}\}$ and $n^{\text{wall}}$ is the number of the wall segment. The affordance of a wall is passively provided and is defined as $\mathbf{a}^{\text{wall}}_{i}$ with the model shown in Sec. 2.3. The passive affordance vector is the normal vector of $\mathbf{s}^{\text{wall}}_{i}$ located in the middle with the direction pointing towards the constrained area. Its magnitude is scaled to half of $\mathbf{s}^{\text{wall}}_{i}$ as the manipulandum is generally not receiving any affordance from a wall segment based on our visual affordance model. The moving direction for the manipulandum to the exit can be obtained by the following equation:
@@
We define “repositioning” as moving the tool closer to the object and realigning $\mathbf{p}_{\ast}$ with $\mathbf{p}^{\text{obj}}$ by $k$ amount. The value of $k$ is determined empirically. In “rotation-dragging”, the tool maintains contact with the manipulandum when it rotates by a certain angle as $\angle{\text{rot}}$ shown in Fig. 5(b) and moves outwards by extending $\overrightarrow{\mathbf{p}^{\text{ee}}_{\tau}\mathbf{p}^{\text{ee}}_{\text{rot}}}$ by a $w>0$ amount.
@@ 3 Results
To validate our methodology in terms of accuracy and robustness, we have conducted around 200 experiments in a dual-arm robot system. In the experiment, two sets of UR-3 robotic arms are used and GPT 4o-mini is implemented for task decomposition. Three types of tools are selected which are a stick, an L-shaped hook, and a Y-shaped hook (see Fig.1). Different tool combinations are evaluated with diverse movement directions and tasks. Various masses of the manipulandum are tested and validated. Since this is a vision-based controller, the mass of the manipulandum does not significantly affect the results. Therefore, the manipulandum’s mass is omitted from this section. A RealSense D415 captures the images of the whole process. Data is passed to a Linux-based computer with the Robot Operating System (ROS) for image process and robot control. Aruco code is used for providing accurate pose tracing in real time. The average inference time is approximately 0.158 seconds for tool analysis and around 1.51 seconds for LLM processing. Since these operations are completed before the robot begins its movements, their effect on overall system responsiveness remains minimal.
@@
3.1 Single-Arm Robot with a Single Tool
We first evaluate the task decomposition performance of LLM. For that, a tool and a blue manipulandum are placed on the table with the target given as shown in Fig. 6. The task is to manipulate the manipulandum within a close distance, which is sufficient for a single-arm robot. The embedded information which contains the task, the environment, and the geometry of the tool is passed to the LLM. In the experiment shown in Fig. 6(a), the robot executes the subtasks generated by the high-level symbolic task planner which include: grasp(right, hook); approach(right, hook, block); interact(right, hook, block, target); release(right, hook). where the right arm first moves and grasps the hook, then moves the block to the target, and lastly releases the tool back to its original place. In a non-single tool scenario, where two tools are available on the desk as shown in Fig. 6(c), the task planner selects the nearest tool based on the embedded information to push the block towards the target. The experiment showcases the application of the proposed affordance and manoeuvrability model in locating the highest manoeuvrability region for manipulandum transportation. During the manipulation stage, the manipulandum is kept within the highest manoeuvrability region (indicated with a red circle in Fig. 6) to receive affordance effectively from the tool. The minimisation of the error between the $\mathbf{p}^{\text{obj}}$ and the $\mathbf{p}^{\text{target}}$ for each experiment is shown in Fig. 7. These results corroborate that the proposed method can be used to actively drive a robot to manipulate an object via a tool.
@@ 3.2 Dual-Arm Robot with Long-Horizon Task
The long-horizon task performance is evaluated with the tool-sharing ability. Assuming there is only one tool available, it has to be shared among the dual-arm robot. Fig. 8(c) demonstrates the tool is passed to another arm once the manipulandum is pushed to the middle of the table. The manipulandum is moved accurately to the target with motion-decomposed: $\mathtt{grasp};$ $\mathtt{approach};$ $\mathtt{interact};$ $\mathtt{pass};$ $\mathtt{release};$ $\mathtt{grasp};$ $\mathtt{approach};$ $\mathtt{interact};$ $\mathtt{release}$’ where the left arm releases the tool once it is done and the right picks up the tool to continue moving the manipulandum. Though the hook is in a two-link geometry, the pushing is afforded by the right side of the tool (a single segment) with the highest manoeuvrability region.

### v6

@@ Abstract
firstofonefirstofonefirstofone
t1t1footnotetext: This work is supported in part by the Research Grants Council of Hong Kong under grant C4042-23GF and in part by the National Natural Science Foundation of China (NSFC) under Grant No. 62403211.
@@ 1.1 Related Works
Transitioning from the critical role of task planning, it is evident that effective tool use is inherently tied to understanding the relationship between tools and objects [38]. Indeed, the success of a given tool-object manipulation task largely depends on the appropriate selection of the tool, which necessitates a nuanced comprehension of how tools interact with various objects in their environment. For example, robots can identify the tool type, potential uses, and contact approaches based on the tool’s geometry, see e.g., [28, 30]. In [41], tool features are learned through observation of the task’s effects and experimental validation of feature hypotheses. Affordance models are a common technique used for tool feature selection [3, 31, 44] and tool classification [37, 44, 34]. The relation between tool actions and their effects on objects is explored in [6, 34], where robots acquire affordance knowledge through predefined actions (e.g., pull, push, rotate). Recently, researchers have also explored the use of LLM in accelerating affordance learning in tool manipulation [30]. Some works have studied tool-based manipulation under constraints and from demonstrations [5]. Non-prehensile object manipulation strategies have been used in [10, 32].
@@ 1.2 Contributions
To address this research gap, we propose a novel LLM-based manoeuvrability-driven method with the following key contributions: (1) We develop a geometric-mechanical model that explicitly captures the interaction between tools and objects, enabling accurate representation of their manoeuvrability in various manipulation scenarios; (2) We introduce a non-prehensile manipulation strategy tailored for tools, allowing efficient object manipulation under various spatial and physical constraints without the need for grasping; (3) We conduct real-world experiments on a dual-arm robotic system, validating the proposed methodology through performance evaluations and demonstrating its practical applicability in dynamic environments.
Our approach uniquely integrates LLMs to enhance tool-object interactions, enabling robots to interpret and perform complex non-prehensile tasks through natural language instructions. This integration not only improves dynamic adaptability to different manipulation scenarios but also promotes more intuitive human-robot collaboration, increasing the effectiveness of dual-arm tool-object manipulation.
The rest of the manuscript is organised as follows: Sec. 2 presents the methodology, Sec. 3 presents the results, Sec. 4 discusses advantages, limitations, and gives final conclusions.
@@ 2.1 Problem Formulation
We define a tool as a manipulable object that is graspable by a robot, a manipulandum [28] as an object (e.g. a block) that is manipulated via a tool, and a wall as a static non-manipulable object. Tool use by robots is challenging as the tools can have various shapes, the environment can be dynamic, and the contact between the tool and the manipulandum may be hard to maintain in a long-horizon task. In this study, we focus on using the side part of a tool to interact with the manipulandum. Depending on the geometric features of a tool and a wall, the available affordance for manoeuvring a manipulandum may be different. Affordance here refers to the available action-effects offered by the tool or the environment. In this work, we classify affordance into two types: active and passive. Active affordance is given from a manipulable object, i.e. a tool, and it is directly related to the manoeuvrability when driving a manipulandum. A passive affordance is given by a static non-manipulable object.
@@ 2.2 LLM-Based High-Level Symbolic Task Planner
To obtain a valid task decomposition for a long-horizon task, the system needs to understand the requirements and generate an executable subtask list. We develop a symbolic task planner that takes natural language instructions with scene descriptions as input, and outputs a list of high-level subtasks. The list involves the tool selection/sharing between two arms, the sequence to manipulate the tool with the manipulandum, and the interaction between the two arms. The model is fine-tuned using approximately 20,000 example data lists, specifically tailored for our non-prehensile tool object manipulation scenario. During the fine-tuning stage, we utilized a program to create 20,000 distinct environmental setups by randomly varying the poses of the robot, tool, block, and target within a finite combination space. To ensure data quality and optimality, each generated setup was validated using rule-based filters that enforced logical and task-relevant constraints, ensuring that only feasible and meaningful manipulation scenarios were retained. This strategy produced a dataset covering both common and rare configurations, enabling the LLM to learn robust mappings between scene layouts and corresponding task plans. By framing task decomposition as a classification problem, the LLM can effectively associate each setup with a specific list of motion functions. This design enhances its ability to generate consistent, physically grounded predictions and reduces the likelihood of producing hallucinated or infeasible task sequences.
@@
The LLM interprets the output of $f(\mathbf{L},\mathbf{o})$ to generate a subtask list $\{\mathbf{l}_{1},\mathbf{l}_{2},\dots\}\leftarrow f_{\text{llm}}(f(\mathbf{L},\mathbf{o}))$ where $\mathbf{l}_{i}$ is a subtask describing the manipulation phase of each robot and corresponds to a high-level robot motion function. The motion functions are designed to be simple and specify a short-term goal of the concerned object (these functions omit low-level motion commands). For simplicity, here we use m to represent the manipulandum in the following function definitions. We use grasp(arm, tool) for grasping a tool with the robot arm; approach(arm, tool, m) for approaching the location of m with tool using arm; interact(arm, tool, m, goal) for moving m to the goal location with the tool; stepping(arm, tool, m) for moving m out from the bounded area with the tool of the arm through contact pulsing motions; pass(arm1, tool, m, arm2) for passing m to another arm’s workspace; release(arm, tool) for releasing the tool back to its original place with the arm.
@@ 2.4 Manoeuvrability Analysis
A tool can push the manipulandum from the side, from the tip, or from other areas. However, the relative location of the manipulandum with respect to the tool affects its manoeuvrability. In other words, the affordance provided by the tool is proportional to manoeuvrability. Consider using a rotating stick to push an object with its end tip. In this situation, the tool may lose contact with the manipulandum as it rolls outwards; hence, the manoeuvrability of this point is low. On the other hand, the midpoint of the stick has a high manoeuvrability, which proportionally decreases as the contact point is further away from the midpoint. This behaviour can be modelled with a Gaussian function, where its centre is the segment’s centre and the peak height is half the segment’s length, see Fig. 3(b). We refer to this region as an affordance area.
@@
Tool regions with high values in the image $\mathbf{M}$ reflect a high manoeuvrability. These computed manoeuvrability values are useful to determine the location where the tool interacts with the manipulandum. To determine the centre of the object, we then expand the contour of the tool by the object’s radius $r^{\text{obj}}$. This contour is downsampled with the Ramer-Douglas-Peucker algorithm, then parameterised with the spline fitting technique. To extract key features of the tool geometry, we use a sliding window strategy to examine a small number of neighbouring points. Let $\mathcal{C}$ be the contour of the tool expanded by $r^{\text{obj}}$. The key features of the tool geometry are extracted using the following equation:
@@
To compute the minimal number of key points (denoted as $\mathbf{P}^{\text{key}}=\{\mathbf{p}^{\text{key}}_{1},\mathbf{p}^{\text{key}}_{2},\dots\}$) that capture the highest manoeuvrability among feature points, we use the density‐based clustering algorithm. By integrating the affordance areas we obtained earlier, we can filter out some redundant key points. For example, if there exists a point $\mathbf{p}^{\text{key}}_{i}$ located outside the affordance area (visualised in Fig. 3(b)), we consider this point as redundant. All the non-redundant points are then grouped into $\mathbf{P}^{\diamond}=\{\mathbf{p}^{\diamond}_{1},\mathbf{p}^{\diamond}_{2},\dots\}$. To find the point in $\mathbf{P}^{\diamond}$ with the highest manoeuvrability (defined as $\mathbf{p}_{\ast}$), we use the manoeuvrability matrix $\mathbf{M}$ and distance between $\mathbf{p}^{\diamond}_{i}$ and $\mathbf{a}_{\ast}$ as described in the metric below:
@@ 2.5.2 Motion Strategy
where the coordinates of $\mathbf{p}^{\text{tool}}$ can be determined by finding a point $\mathbf{p}=(x,y)$ where it minimizes the distance between $(\mathbf{p}_{\ast},\mathbf{p}^{\text{obj}})$ with $f(\mathbf{p})$ and $(\mathbf{p}^{\text{tool}},\mathbf{p}^{\text{int}}_{\text{start}})$; then translates along the $x$ and $y$ axes until it reaches $\mathbf{p}^{\text{int,x,y}}_{\text{end}}$ with $k_{\text{int}}(\mathbf{p}^{\text{int,x,y}}_{\text{end}}-\mathbf{p}^{\text{tool}})$, where $k_{\text{int}}$ is determined empirically; lastly, the tool is rotated to align with the orientation of $\mathbf{p}^{\text{int}}_{\text{end}}$.
@@ 2.6 Application with Environmental Constraints
Consider the manipulandum is tightly confined within a concave-shaped wall, as shown in Fig. 5(a), with an unknown exit and assume that the tool can enter the constrained area. To move the manipulandum out of the bounded area with a small movement space, we determine the direction from the manipulandum to the exit by considering the overall affordance of the wall boundary. We denote this direction vector as $\mathbf{v}^{\text{exit}}$, and its magnitude is defined as the minimum travel distance for the manipulandum. Consider the inner edge of the wall as a segment $\mathbf{s}^{\text{wall}}_{i}$ where $i=\{1,\dots,n^{\text{wall}}\}$ and $n^{\text{wall}}$ is the number of the wall segment. The affordance of a wall is passively provided and is defined as $\mathbf{a}^{\text{wall}}_{i}$ with the model shown in Sec. 2.3. The passive affordance vector is the normal vector of $\mathbf{s}^{\text{wall}}_{i}$ located in the middle with the direction pointing towards the constrained area. Its magnitude is scaled to half of $\mathbf{s}^{\text{wall}}_{i}$ as the manipulandum is generally not receiving any affordance from a wall segment based on our visual affordance model. The moving direction for the manipulandum to the exit can be obtained by the following equation:
@@
We define “repositioning” as moving the tool closer to the object and realigning $\mathbf{p}_{\ast}$ with $\mathbf{p}^{\text{obj}}$ by $k$ amount. The value of $k$ is determined empirically, representing the spatial offset between the tool and the object. A larger $k$ allows a wider clearance before contact, while a smaller $k$ brings the tool closer, increasing precision but also the risk of collision. In “rotation-dragging”, the tool maintains contact with the manipulandum when it rotates by a certain angle as $\angle{\text{rot}}$ shown in Fig. 5(b) and moves outwards by extending $\overrightarrow{\mathbf{p}^{\text{ee}}_{\tau}\mathbf{p}^{\text{ee}}_{\text{rot}}}$ by a $w>0$ amount. If $\angle{\text{rot}}$ or $w$ are excessively large, the tool may jam or cause damage in the constrained area; conversely, values that are too small reduce efficiency by requiring more iterations to complete the manipulation. It is a trade-off between maintaining stability and achieving motion efficiency.
@@ 3 Results
To evaluate the proposed framework in terms of accuracy, robustness, and practical feasibility, approximately 200 experiments are conducted using a dual-arm UR-3 robotic system. The fine-tuning of the large language model (GPT-4o-mini) is performed in the cloud on GPU-enabled servers, and during deployment, the robotic system accesses the trained model through a secure API connection for inference. Three types of tools are selected, which are a stick, an L-shaped hook, and a Y-shaped hook (see Fig.1). Different combinations of these tools were evaluated under diverse movement directions and task objectives. Various masses of the manipulandum are tested, but due to the minimal impact on the vision-based controller, mass is excluded from this section. The experimental tasks covered a wide range of scenarios, including close-range manipulation with single and multiple tools, long-horizon (single and tool-sharing) operations, and manipulation within constrained environments. An Intel RealSense D415 captures the images of the whole process. Data is passed to a Linux-based computer with the Robot Operating System (ROS) for image processing and robot control. Aruco markers are used for providing accurate pose tracing in real time. The average inference time is approximately 0.158 seconds for tool analysis and around 1.51 seconds for LLM processing. Since these operations are completed prior to robot movement, their latency had minimal influence on overall system responsiveness.
@@
3.1 Single-Arm Robot
We first evaluate the task decomposition performance of LLM. For that, a tool and a blue manipulandum are placed on the table with the target given as shown in Fig. 6. The task is to manipulate the manipulandum within a close distance, which is sufficient for a single-arm robot. The embedded information, which contains the task, the environment, and the geometry of the tool, is passed to the LLM. In the experiment shown in Fig. 6(a), the robot executes the subtasks generated by the high-level symbolic task planner, which include: grasp(right, hook); approach(right, hook, block); interact(right, hook, block, target); release(right, hook). The right arm first moves and grasps the hook, then moves the block to the target, and lastly releases the tool back to its original place.
In a non-single tool scenario, where two tools are available on the desk as shown in Fig. 6(c), the task planner selects the nearest tool based on the embedded information to push the block towards the target. The experiment showcases the application of the proposed affordance and manoeuvrability model in locating the highest manoeuvrability region for manipulandum transportation. During the manipulation stage, the manipulandum is kept within the highest manoeuvrability region (indicated with a red circle in Fig. 6) to receive affordance effectively from the tool. The minimisation of the error between the $\mathbf{p}^{\text{obj}}$ and the $\mathbf{p}^{\text{target}}$ for each experiment is shown in Fig. 7. These results corroborate that the proposed method can be used to actively drive a robot to manipulate an object via a tool.
@@ 3.2 Dual-Arm Robot with Long-Horizon Task
Table 1: Manipulation Accuracy Across Subtasks for Different Tools
Subtask Stick Hook Y-hook
Grasp 100% 100% 100%
Approach 96% 97% 95%
Interact 91% 92% 92%
Pass 92% 93% 93%
Release 100% 100% 100%
Fig. 10: Comparison of success rate under different state-of-the-art methodologies. FT states for fine-tuning, SRST states for a single-arm robot with a single tool, Dual refers to dual arms collaboration with two tools, and Sharing refers to tool-sharing collaboration.
The long-horizon task performance is evaluated with the tool-sharing ability. Assuming there is only one tool available, it has to be shared among the dual-arm robot. Fig. 8(c) demonstrates that the tool is passed to another arm once the manipulandum is pushed to the middle of the table. The manipulandum is moved accurately to the target with motion-decomposed: $\mathtt{grasp};$ $\mathtt{approach};$ $\mathtt{interact};$ $\mathtt{pass};$ $\mathtt{release};$ $\mathtt{grasp};$ $\mathtt{approach};$ $\mathtt{interact};$ $\mathtt{release}$’ where the left arm releases the tool once it is done and the right picks up the tool to continue moving the manipulandum. Though the hook is in a two-link geometry, the pushing is afforded by the right side of the tool (a single segment) with the highest manoeuvrability region.
@@
To further evaluate subtask performance, Table 1 shows the manipulation accuracy, measured as the alignment between $\mathbf{p}^{\text{obj}}$ and the desired point for the approach, interact, and pass subtasks, and the success rate for the grasp and release subtasks across different tools. The results show 100% success for grasp and release, while the rest of the subtasks maintain high accuracy above 90% for all tools. These findings demonstrate that each subtask is executed reliably and that the proposed framework achieves robust performance across diverse tool geometries in long-horizon dual-arm tasks.